OpenStudy (anonymous):

Divide using synthetic division, and write a summary statement in fraction form.

OpenStudy (anonymous):

\[\frac{ 2x^4-x^3-15x^2+3x }{ x+3 }\]

OpenStudy (anonymous):

\[ 2x3 - 7x2 + 6x -15 + \frac{ 45 }{ x+3 }\] \[ 2x3 - 5x2 + 3 +\frac{ -9 }{ x+3 }\] \[2x3 - 7x2 + 6x -15 +\frac{ -45 }{ ?x+3}\] \[2x3 + 5x2 + 3 + \frac{ 9 }{ x+3 }\]